libosanajustine:

Water is being drained through a hole at the bottom of a tank at a velocity of 10 m/s. For a tank diameter of 1.18 m and a hole diameter of 5 cm, find the rate of change of the water level in the tank in m/min.

darkknight:

@darkknight taggin myself for tomorrow, this is a related rates calc problem

jhonyy9:

- this 10 m/s is the first water level change so i think bc. the measures of tank dont changed we can use the rule of 3 simple 10 m change in 1 sec x m change in 60 sec = 1 min. x = 60*10/1 = 60*10 = 600 m/min.

jhonyy9:

@supie any idea here or do you agree me ? @Laylalyssa

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @supie any idea here or do you agree me ? @Laylalyssa \(\color{#0cbb34}{\text{End of Quote}}\) I agree,

jhonyy9:

ty

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ty \(\color{#0cbb34}{\text{End of Quote}}\) yw

YECR:

Did u solve it are you good with it? if ur good then close the question

darkknight:

Water is being drained through a hole at the bottom of a tank at a velocity of 10 m/s. For a tank diameter of 1.18 m and a hole diameter of 5 cm, find the rate of change of the water level in the tank in m/min. convert everything to cm per sec water draining from hole at 1000cm/sec diameter = 118 cm and radius = 59cm the hole has a diameter of 5 cm

darkknight:

u find the rate of change by finding the change of velocity, by using formula for a cylindrical tank \[V=\pi r^2h\] We need to isolate the radius and solve for the height not sure how to do this since you did not give us a height. Unless I'm bonkers and don't understand the problem

darkknight:

@jhonyy9 this feels like a related rates problem (calc) bc i learned it a month ago, I don't think you did this one right. @Laylalyssa any idea

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight @jhonyy9 this feels like a related rates problem (calc) bc i learned it a month ago, I don't think you did this one right. @Laylalyssa any idea \(\color{#0cbb34}{\text{End of Quote}}\) no, bye